# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ 5 Conclusion and Discussion
This paper introduces a unified framework for conducting valid statistical inference using machine learning imputations in datasets characterized by general, multiple different missing data patterns under the MAR assumption. By leveraging the concept of missingness patterns and building upon the theoretical foundations of AIPW estimators, our work significantly extends the applicability of prediction-based inference beyond the more restrictive MCAR assumption and simple missing data structures, such as missing outcomes or blockwise missing covariates. A key practical advantage of our proposed methods is their adaptability to any Z-estimation framework, allowing researchers to utilize existing statistical software for weighted complete-case analysis with minimal additional implementation effort.
@@
Despite these important contributions, our work has several limitations that inform avenues for future research. First, a central assumption of our methodology is the correct specification of the propensity score model, which governs the probability of observing different missingness patterns. This assumption is the price we pay in the challenging scenario where prediction-based inference is most needed: analysts typically have minimal or no control of the prediction models. As such, assumptions on the prediction models need to be avoided, and the correct specification of the propensity score model comes as the price. Although we demonstrate valid inference when propensity scores are known or consistently estimated from a correctly specified model, results from Figure S4 and S5 show that propensity score model misspecification will lead to biased estimation when the ML predictions are not sufficiently precise. In practice, however, correctly specifying the propensity score model, especially when the number of missingness patterns $K$ is large, is challenging if not impossible. As a result, although this paper provide a theoretical foundation for an arbitrary $K$, it is more practically feasible when $K$ is small. Future work should focus on robust propensity score model against misspecification. Furthermore, it is worth noting that results from Figure S4 and S5 exhibits empirical evidence of the double robustness of the proposed PS-PPI method, where PS-PPI achieve the 95% nominal coverage rate when the quality of the ML predictions are precise enough (or perfect). Recalling that the PS-PPI estimator is an AIPW estimator, a thorough theoretical investigation to the potential double robustness property of the PS-PPI estimator is needed as a future work to clarify the quality requirement of the ML predictions to bypass the challenging modeling process of the propensity score models.
Second, another practical limitation of PS-PPI is that it imposes assumptions to the minimum size of data within each missingness patterns to guarantee the learnability of the parameter of interest. For example, to fit a linear regression analysis model with $p$ parameters (including the intercept), at least $p+1$ data points are needed. In this case, PS-PPI requires each missingness pattern within the data to contain at least $p+1$ data points, and if such requirement is not satisfied, then data with that missingness pattern have to be dropped from the dataset.
Third, while PS-PPI provides a data-driven approach for selecting tuning parameters ($\mathbf{W}_{k}$) that guarantees efficiency gains, the assumption of identical $\mathbf{W}_{k}$ across strata was made for simplicity and ease of implementation. Future work could explore more adaptive stratum-specific weight selection to potentially achieve even greater efficiency, although this may involve more complex optimization procedures. Furthermore, the number of possible missingness patterns can grow exponentially with the number of variables. Although in many practical applications, the number of observed patterns is manageable, scenarios with a very large number of distinct patterns could pose computational challenges for the stratification and subsequent analyses. Research into methods for adaptively grouping or regularizing coarsening patterns could be beneficial.
Fourth, while our framework is versatile within Z-estimation problems, its extension to other inferential paradigms or specific complex data types, such as longitudinal or spatial data where the temporal and spatial dependency of missingness is critical, or high-dimensional settings where variable selection and inference are performed simultaneously, presents fertile ground for future research. The current work also focuses on the MAR assumption. Extending these prediction-powered inference ideas to handle Missing Not At Random (MNAR) scenarios, while notoriously challenging due to identifiability issues, would be a significant theoretical and practical advancement.
The broader implications of this research are multifold. By providing a more general and statistically sound framework for using ML-imputed data, PS-PPI can encourage the adoption of advanced predictive modeling in substantive research areas without sacrificing inferential rigor. This can lead to more powerful and nuanced analyses of existing datasets, potentially uncovering new scientific insights that were previously obscured by missing information or by the limitations of traditional missing data techniques. Moreover, the ability to validly incorporate ML predictions could influence how researchers approach data collection, potentially allowing for more targeted and cost-effective acquisition of complete data on subsets of variables, knowing that robust inferential methods exist to handle the resulting missingness.
@@
$\widehat{{\bm{\theta}}}_{\text{PP}}$ An estimator from outcome-based prediction-powered inference (Angelopoulos et al., 2023a, ).
@@
$\widehat{{\bm{\theta}}}_{\text{Chen}}$ An outcome-based PPI++ style estimator from Gronsbell et al., (2024); Chen and Chen, (2000) that guarantees to be more efficient than $\widehat{{\bm{\theta}}}_{\text{CC}}$.
$\widehat{{\bm{\theta}}}_{\text{PTD}}$ An estimator from Kluger et al., (2025) for blockwise covariates missing with nonuniform sampling weights.
@@
$\widehat{\Sigma}_{{\bm{\theta}}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$.
$\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ The asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$.
$\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ The asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, $k,k^{\prime}\in\{1,...,K\}$. When $k=k^{\prime}$, it becomes the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$.
$\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ The asymptotic variance of $\widehat{{\bm{\gamma}}}_{2,k}$.
$\widehat{\Sigma}_{\text{PS-PPI}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
@@
We first stack the estimating equations and obtain:
@@
By working on the stacked estimating equations we have the following asymptotic linear expansion:
$\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{(i)}+o_{p}(1),$
@@
$\dot{\bm{\Psi}}=\mathbb{E}\Bigl{[}\tfrac{\partial\bm{\Psi}}{\partial\bm{\vartheta}}\Bigr{|}_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\Bigr{]}=\mathbb{E}\!\left[\,\begin{bmatrix}\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\omega}}}\\[6.0pt] \frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}_{1,1}}{\partial{\bm{\omega}}}\\[2.0pt] \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\[2.0pt] \frac{\partial{\bm{m}}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{m}}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{m}}}{\partial{\bm{\omega}}}\end{bmatrix}_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\,\right].$
From van der Vaart, (Theorem 5.21 2000) and Assumption 2, $\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})$ is asymptotically normal. To show that $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal, we first introduce the following notation:
@@
$\dot{{\bm{\psi}}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\theta}}}\Bigr{|}_{{\bm{\theta}}={\bm{\theta}}^{\ast}}\right],\leavevmode\nobreak\ \dot{{\bm{\psi}}}_{k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{1,k}}{\partial{\bm{\gamma}}_{1,k}}\Bigr{|}_{{\bm{\gamma}}_{1,k}={\bm{\gamma}}^{\ast}_{k}}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{2,k}}{\partial{\bm{\gamma}}_{2,k}}\Bigr{|}_{{\bm{\gamma}}_{2,k}={\bm{\gamma}}^{\ast}_{k}}\right],$
@@
$\dot{{\bm{\psi}}}^{\bm{\omega}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{\text{WCC}}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right],\leavevmode\nobreak\ \dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{1,k}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right],\leavevmode\nobreak\ \dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}_{2,k}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right],\leavevmode\nobreak\ \mathbf{M}=\mathbb{E}\left[\frac{\partial{\bm{m}}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right].$
@@
Then, by noticing that all the non-diagonal elements except those in the last row are $\bm{0}$, we have:
@@
$\dot{\bm{\Psi}}=\begin{bmatrix}\dot{{\bm{\psi}}}&\bm{0}&\cdots&\dot{{\bm{\psi}}}^{{\bm{\omega}}}\\ \bm{0}&\dot{{\bm{\psi}}}_{1}&\cdots&\dot{{\bm{\psi}}}^{{\bm{\omega}}}_{1,1}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&\mathbf{M}\end{bmatrix}=\begin{bmatrix}\mathbf{J}_{1}&\mathbf{J}_{2}\\ \bm{0}&\mathbf{M}\end{bmatrix}$
@@
where $\mathbf{J}_{1}=\text{diag}(\dot{{\bm{\psi}}},\dot{{\bm{\psi}}}_{1},\cdots,\dot{{\bm{\psi}}}_{K},\dot{{\bm{\psi}}}_{1},\cdots\dot{{\bm{\psi}}}_{K})$, $\mathbf{J}_{2}=(\dot{{\bm{\psi}}}^{\bm{\omega}},\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,1},\cdots,\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,K},\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,1},\cdots\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,K})^{\top}$. Hence, from Lu and Shiou, (2002) we can obtain the inversion of the upper triangle $2\times 2$ block matrix $\dot{\bm{\Psi}}$:
@@
$\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}^{-1}&\bm{0}&\cdots&-\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[6.0pt] \bm{0}&\dot{{\bm{\psi}}}_{1}^{-1}&\cdots&-\dot{{\bm{\psi}}}_{1}^{-1}\dot{{\bm{\psi}}}_{1,1}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[2.0pt] \vdots&\vdots&\ddots&\vdots\\[2.0pt] \bm{0}&\bm{0}&\cdots&\mathbf{M}^{-1}\end{bmatrix}$

### v2

@@ 6 Conclusion and Discussion
This paper introduces a unified framework for conducting valid statistical inference using machine learning imputations in datasets characterized by general, multiple different missing data patterns under the MAR assumption. By leveraging the concept of missingness patterns and building upon the theoretical foundations of AIPW estimators, our work significantly extends the applicability of prediction-based inference beyond the more restrictive MCAR assumption and simple missing data structures, such as missing outcomes or missing multiple covariates. A key practical advantage of our proposed methods is their adaptability to any Z-estimation framework, allowing researchers to utilize existing statistical software for weighted complete-case analysis with minimal additional implementation effort. The method introduced here can integrate individuals across diverse missingness patterns, hence maximizing data utility and inclusiveness while ensuring accuracy.
@@
Despite these important contributions, our work has several limitations that inform avenues for future research. First, a central assumption of our methodology is the correct specification of the propensity score model, which governs the probability of observing different missingness patterns. This assumption is the price we pay in the challenging scenario where prediction-based inference is most needed: analysts typically have minimal or no control over the prediction models. As such, assumptions on the prediction models need to be avoided, and the correct specification of the propensity score model comes at a price. Although we demonstrate valid inference when propensity scores are known or consistently estimated from a correctly specified model, results from Figure S4 and S5 in the Supplementary Materials show that propensity score model misspecification will lead to biased estimation when the ML predictions are not sufficiently precise. It is worth noting that results from Figure S4 and S5 in the Supplementary Materials show that PS-PPI achieves the 95% nominal coverage rate when the ML predictions are precise enough (or perfect). Future work should focus on mitigating the effect of propensity score model misspecification. Second, another practical limitation of PS-PPI is that it imposes assumptions on the minimum size of data within each missingness pattern to guarantee the learnability of the parameter of interest. For example, to fit a linear regression analysis model with $p$ parameters (including the intercept), at least $p+1$ data points are needed. In this case, PS-PPI requires each missingness pattern within the data to contain at least $p+1$ data points, and if this requirement is not satisfied, then data with that missingness pattern have to be dropped from the dataset. Third, while PS-PPI provides a data-driven approach for selecting tuning parameters ($\mathbf{W}_{k}$) that guarantees efficiency gains, the assumption of identical $\mathbf{W}_{k}$ across strata was made for simplicity and ease of implementation. Future work could explore more adaptive stratum-specific weight selection to potentially achieve even greater efficiency, although this may involve more complex optimization procedures. Fourth, the number of possible missingness patterns can grow exponentially with the number of variables. Although in many practical applications, the number of observed patterns is manageable, scenarios with a very large number of distinct patterns could pose computational challenges. Future work can improve the utility of the framework by developing adaptive procedures to use data from a small number of the most informative missingness patterns. Fifth, while our framework is versatile within Z-estimation problems, its extension to other inferential paradigms or specific complex data types, such as longitudinal or spatial data where the temporal and spatial dependency of missingness is critical, or high-dimensional settings where variable selection and inference are performed simultaneously, presents fertile ground for future research. The current work also focuses on the MAR assumption. Extending these prediction-powered inference ideas to handle missing not at random (MNAR) scenarios, while notoriously challenging due to identifiability issues, would be a significant theoretical and practical advancement. We leave these directions for future work.
Acknowledgment
We gratefully acknowledge All of Us participants for their contributions, without whom this research would not have been possible. We also thank the National Institutes of Health’s All of Us Research Program for making available the participant data examined in this study. The authors acknowledge the use of Large language models (LLM)-assisted technologies by GPT-5 to refine the language and grammar. All LLM suggestions were evaluated by the first and the last authors prior to their incorporation into the manuscript. LLMs were strictly limited to proposing edits to the initial drafts and were not used to generate any new materials.
Supplementary Materials
Supplementary Materials contain a glossary of notation, proofs, supporting figures, longer derivations, and additional results. Code to reproduce simulations and data analysis will be made available online at https://github.com/chenxran/ps-ppi.
@@ Acknowledgment
$\widehat{{\bm{\theta}}}_{\text{Full}}$ An estimator obtained from the fully-observed data.
@@
$\widehat{{\bm{\theta}}}_{\text{PP}}$ An estimator from outcome-based prediction-powered inference (angelopoulos2023prediction).
@@
$\widehat{{\bm{\theta}}}_{\text{Chen}}$ An outcome-based PPI++ style estimator from gronsbell2024another; chen2000unified that guarantees to be more efficient than $\widehat{{\bm{\theta}}}_{\text{CC}}$.
$\widehat{{\bm{\theta}}}_{\text{PTD}}$ An estimator from kluger2025prediction for missing multiple covariates with nonuniform sampling weights.
@@
${\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})$ The estimating equation for $\widehat{{\bm{\theta}}}_{\text{Full}}$ with the $i$-th data plugged in. It is a function of $\bm{\theta}$} and can be abbreviated as ${\bm{\psi}}^{(i)}$. When the true parameter $\bm{\theta}$$^{\ast}$ is plugged in, i.e., ${\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$, it can be abbreviated as ${\bm{\psi}}^{\ast(i)}$.
${\bm{\psi}}^{(i)}_{\text{WCC}}$ The estimating equation for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ with the $i$-th data point plugged in.
${\bm{\psi}}^{(i)}_{1,k}$ The estimating equation for $\widehat{{\bm{\gamma}}}_{1,k}$ with the $i$-th data point plugged in.
${\bm{\psi}}^{(i)}_{2,k}$ The estimating equation for $\widehat{{\bm{\gamma}}}_{2,k}$ with the $i$-th data point plugged in.
${\bm{m}}^{(i)}$ The estimating equation of the propensity score model with the $i$-th data point plugged in.
$\Sigma_{{\bm{\theta}}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{WCC}}$.
$\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ The asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$.
$\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ The asymptotic covariance between $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, $k,k^{\prime}\in\{1,...,K\}$. When $k=k^{\prime}$, it becomes the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$.
$\Sigma_{{\bm{\gamma}}_{2,k}}$ The asymptotic variance of $\widehat{{\bm{\gamma}}}_{2,k}$.
$\Sigma_{\text{PS-PPI}}$ The asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
@@
To prove Theorem 1, we first stack the estimating equations and obtain:
@@
The proof sketch of the Theorem 1 is that: 1) we derive the asymptotic linear expansion of the stacked estimator $\widehat{\bm{\vartheta}}$; 2) we show that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is a linear transformation of the elements in $\widehat{\bm{\vartheta}}$, thereby obtaining the asymptotic linear expansion of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$.
@@
By working on the stacked estimating equations, we have the following asymptotic linear expansion:
$\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{\ast(i)}+o_{p}(1),$
@@
$\dot{\bm{\Psi}}=\mathbb{E}\Bigl[\tfrac{\partial\bm{\Psi}^{(i)}}{\partial\bm{\vartheta}}\Bigr|_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\Bigr]=\mathbb{E}\!\left[\,\begin{bmatrix}\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\omega}}}\\[6.0pt] \frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{\psi}}^{(i)}_{1,1}}{\partial{\bm{\omega}}}\\[2.0pt] \vdots&\vdots&\ddots&\vdots&\vdots&\ddots&\vdots&\vdots\\[2.0pt] \frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\theta}}}&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{1,1}}&\cdots&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{1,K}}&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{2,1}}&\cdots&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\gamma}}_{2,K}}&\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\omega}}}\end{bmatrix}_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\,\right].$
To obtain the inverse of $\dot{\bm{\Psi}}$, we first introduce the following notation:
@@
$\dot{{\bm{\psi}}}_{\infty}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\theta}}}\Bigr|_{({\bm{\theta}},{\bm{\omega}})=({\bm{\theta}}^{\ast},{\bm{\omega}}^{\ast})}\right],\penalty 10000\ \dot{{\bm{\psi}}}_{k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\gamma}}_{1,k}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\gamma}}_{2,k}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right],$
@@
$\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{\text{WCC}}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\theta}},{\bm{\omega}})=({\bm{\theta}}^{\ast},{\bm{\omega}}^{\ast})}\right],\penalty 10000\ \dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right],$
$\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right],\penalty 10000\ \mathbf{M}=\mathbb{E}\left[\frac{\partial{\bm{m}}^{(i)}}{\partial{\bm{\omega}}}\Bigr|_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right].$
Note that according to Proposition 1, $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\gamma}}_{1,k}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]=\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\gamma}}_{2,k}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]$, therefore we are able to use a single notation to represent both. However, when taking derivative on $\bm{\omega}$} instead of ${\bm{\gamma}}_{1,k}$ or ${\bm{\gamma}}_{2,k}$, there is no guarantee that $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{1,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{1,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]$ and $\mathbb{E}\left[\frac{\partial{\bm{\psi}}^{(i)}_{2,k}}{\partial{\bm{\omega}}}\Bigr|_{({\bm{\gamma}}_{2,k},{\bm{\omega}})=({\bm{\gamma}}^{\ast}_{k},{\bm{\omega}}^{\ast})}\right]$ are the same. As a result, we use two different notation to represent the two.
@@
Then, by noticing that all the non-diagonal elements except those in the last column are $\bm{0}$, we have:
@@
$\dot{\bm{\Psi}}=\begin{bmatrix}\dot{{\bm{\psi}}}_{\infty}&\bm{0}&\cdots&\dot{{\bm{\psi}}}_{\infty}^{{\bm{\omega}}}\\ \bm{0}&\dot{{\bm{\psi}}}_{1}&\cdots&\dot{{\bm{\psi}}}^{{\bm{\omega}}}_{1,1}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{0}&\bm{0}&\cdots&\mathbf{M}\end{bmatrix}=\begin{bmatrix}\mathbf{J}_{1}&\mathbf{J}_{2}\\ \bm{0}&\mathbf{M}\end{bmatrix}$
@@
where $\mathbf{J}_{1}=\text{diag}(\dot{{\bm{\psi}}}_{\infty},\dot{{\bm{\psi}}}_{1},\cdots,\dot{{\bm{\psi}}}_{K},\dot{{\bm{\psi}}}_{1},\cdots\dot{{\bm{\psi}}}_{K})$, $\mathbf{J}_{2}=(\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}},\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,1},\cdots,\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,K},\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,1},\cdots\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,K})^{\top}$. From lu2002inverses, the inverse of such a upper triangle $2\times 2$ block matrix $\dot{\bm{\Psi}}$ is:
@@
$\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}^{-1}_{\infty}&\bm{0}&\cdots&-\dot{{\bm{\psi}}}_{\infty}^{-1}\dot{{\bm{\psi}}}_{\infty}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[6.0pt] \bm{0}&\dot{{\bm{\psi}}}_{1}^{-1}&\cdots&-\dot{{\bm{\psi}}}_{1}^{-1}\dot{{\bm{\psi}}}_{1,1}^{{\bm{\omega}}}\mathbf{M}^{-1}\\[2.0pt] \vdots&\vdots&\ddots&\vdots\\[2.0pt] \bm{0}&\bm{0}&\cdots&\mathbf{M}^{-1}\end{bmatrix}$